# MomentSeeker: A Comprehensive Benchmark and A Strong Baseline For Moment Retrieval Within Long Videos

### v2

@@ Abstract
Retrieval augmented generation (RAG) holds great promise in addressing challenges associated with long video understanding. These methods retrieve useful moments from long videos for their presented tasks, thereby enabling multimodal large language models (MLLMs) to generate high-quality answers in a cost-effective way. In this work, we present MomentSeeker, a comprehensive benchmark to evaluate retrieval models’ performance in handling general long-video moment retrieval (LVMR) tasks. MomentSeeker offers three key advantages. First, it incorporates long videos of over 500 seconds on average, making it the first benchmark specialized for long-video moment retrieval. Second, it covers a wide range of task categories (including Moment Search, Caption Alignment, Image-conditioned Moment Search, and Video-conditioned Moment Search) and diverse application scenarios (e.g., sports, movies, cartoons, and ego), making it a comprehensive tool for assessing retrieval models’ general LVMR performance. Additionally, the evaluation tasks are carefully curated through human annotation, ensuring the reliability of assessment. We further fine-tune an MLLM-based LVMR retriever on synthetic data, which demonstrates strong performance on our benchmark. We perform extensive experiments with various popular multimodal retrievers based on our benchmark, whose results highlight the challenges of LVMR and limitations for existing methods. Our created resources is available at https://github.com/yhy-2000/MomentSeeker.
@@ 1 Introduction
Figure 1: Demonstration of our benchmark. Given a natural language query $\textbf{q}_{\textbf{T}}$ (optionally accompanied by an image $\textbf{q}_{\textbf{I}}$ or video $\textbf{q}_{\textbf{V}}$), models should search the provided video to precisely identify the most relevant clips. We marked the parts that directly answer the query with red circles.
@@
2. We conduct extensive experiments with various popular multimodal retrievers based on our benchmark, whose results highlight the challenges of LVMR and limitations for existing methods.
@@ 2.2 Multi-modal Embedding Models
Furthermore, with the rise of large language models, some works attempt to transform multi-modal large models (MLLMs) into universal embedders to construct more powerful representations, such as E5-V Jiang et al. (2024) and MLLM2Vec Jiang et al. (2025). However, in the video domain, a robust universal embedder has yet to emerge. For this reason, we propose V-Embedder, which outperforms existing video embedders across multiple tasks through synthetic-data-driven training.
@@ 5.1 Annotation Overflow
Table 2: Results on the MomentSeeker benchmark. We report R@1 scores averaged across meta-tasks, with all baselines evaluated in a zero-shot setting. Additionally, we provide mAP@5 scores and present detailed performance breakdowns for both metrics on individual sub-tasks in Appendix 8.1.
Method Backbone # Params Meta-Task Average Score Overall
@@
number of tasks - - 5 5 4 4 18
CLIP-style Video Encoder
MM-Ret Zhou et al. (2024) CLIP-Base 149M 23.2 15.4 10.5 10.5 14.9
MagicLens Zhang et al. (2024) CLIP-Large 428M 9.0 2.4 3.2 2.8 4.4
UniIR Wei et al. (2025) CLIP-Large 428M 25.0 15.2 6.4 0.0 10.9
LanguageBind Zhu et al. (2023) CLIP-Large 428M 39.6 16.4 3.2 0.0 14.8
InternVideo2 Wang et al. (2025) ViT 1B 44.6 18.2 4.8 0.0 16.9
@@
CoVR Ventura et al. (2024) BLIP-Large 588M 25.8 17.4 12.3 12.3 17.1
@@
MLLM2VEC Jiang et al. (2025) LLaVA-1.6 8.4B 6.4 6.2 3.0 3.0 4.7
E5V Jiang et al. (2024) LLaVA-1.6 8.4B 25.8 16.8 6.2 5.2 13.5
V-Embedder InternVideo2-Chat 8B 42.2 20.4 15.0 15.8 23.3
@@ 5.2.1 Baselines
MLLM-based Multi-Modality Encoders. Methods like E5V Jiang et al. (2024) and MLLM2Vec Jiang et al. (2025) fall into this category. They utilize large-scale pre-trained multi-modal models to handle flexible modality compositions (e.g., video + text or text alone) and generating multi-modal embeddings.
@@ 5.2.2 Implementation Details
Following previous works Jiang et al. (2025, 2024), we incorporated instructions (e.g., “Represent the given question in one word.” and “Represent the video in one word.”) to guide the model in generating informative embeddings for queries and candidates. All baselines are reproduced using their official implementations with default settings. Instruction usage remains consistent with each baseline’s original configuration. For image-based baselines, we use the middle frame of the video as input. CoVR Ventura et al. (2024) follows its original configuration with 15 frames at a resolution of 384, while LanguageBind, InternVideo2 Wang et al. (2025), and V-Embedder each use 8 frames at a resolution of 224.
@@ 5.3 Main Results
1. CLIP-style video encoders perform well on text-based retrieval tasks but struggle with multi-modal retrieval. InternVideo2 achieves an R@1 of 44.6 on CA and 18.2 on MS but drops to 4.8 on IMS and 0.0 on VMS. Similarly, LanguageBind, the second-best baseline on CA and MS, achieves 39.6 on CA and 16.4 on MS but also falls to 3.2 on IMS 0.0 on VMS. This suggests that despite their strong temporal modeling, these models fail to establish effective multi-modal combination and alignment.
2. Composed Video Retriever excels in multi-modal tasks but struggles with text-based retrieval robustness. CoVR achieves a score of 12.3 on the VMS benchmark but lags significantly in CA performance, scoring 25.8 compared to InternVideo2’s 44.6. This highlights its suboptimal balance between compositional reasoning and general retrieval capabilities.
3. Composed image retrievers demonstrate unexpected generalization to video tasks. MM-Ret scores 10.5 on both IMS and VMS benchmarks, outperforming specialized video encoders like LanguageBind, which scores only 3.2 on IMS. However, its reliance on single-frame inputs restricts temporal understanding, resulting in an overall subpar performance of 17.1.
4. MLLM-based embedding models dominate instruction-based tasks. The MLLM-powered E5V model surpasses video encoder LanguageBind on MS and VMS tasks, achieving 16.8 and 5.2 respectively against LanguageBind’s 16.4 and 0.0. However, E5V trails substantially in CA performance with 22.3 versus LanguageBind’s 39.6, underscoring persistent challenges in fine-grained temporal alignment.
5. V-Embedder achieves top performance across all tasks, demonstrating strong generalization in both traditional retrieval (CA) and complex instructional scenarios (MS, IMS, VMS). Compared to video foundation models, our model performs comparably to InternVideo2 on the traditional CA task. In tasks requiring complex instruction understanding, our model significantly surpasses these baselines, outperforming the strongest existing baseline by 2.2 points on MS and by over 10 points on IMS and VMS, leading to an overall 6.4 points advantage over InternVideo2. This highlights that our MLLM-Embedder effectively balances temporal modeling, cross-modal alignment, and compositional reasoning, making it a strong and general model.

### v3

@@ Abstract
Retrieval augmented generation (RAG) holds great promise in addressing challenges associated with long video understanding. These methods retrieve useful moments from long videos for their presented tasks, thereby enabling multi-modal large language models (MLLMs) to generate high-quality answers in a cost-effective way. In this work, we present MomentSeeker, a comprehensive benchmark to evaluate retrieval models’ performance in handling general long-video moment retrieval (LVMR) tasks. MomentSeeker offers three key advantages. First, it incorporates long videos of over 500 seconds on average, making it the first benchmark specialized for long-video moment retrieval. Second, it covers a wide range of task categories (including Moment Search, Caption Alignment, Image-conditioned Moment Search, and Video-conditioned Moment Search) and diverse application scenarios (e.g., sports, movies, cartoons, and ego), making it a comprehensive tool for assessing retrieval models’ general LVMR performance. Additionally, the evaluation tasks are carefully curated through human annotation, ensuring the reliability of assessment. We further fine-tune an MLLM-based LVMR retriever on synthetic data, which demonstrates strong performance on our benchmark. We perform extensive experiments with various popular multi-modal retrievers based on our benchmark, whose results highlight the challenges of LVMR and limitations for existing methods. Our created resources is available at https://github.com/yhy-2000/MomentSeeker.
@@ 1 Introduction
Figure 1: Demonstration of our benchmark. Given a natural language query $\textbf{q}_{\textbf{T}}$ (optionally accompanied by an image $\textbf{q}_{\textbf{I}}$ or video $\textbf{q}_{\textbf{V}}$), models should search the provided video to precisely identify the most relevant clips. The ground truth clip is pointed by arrows, and we marked the parts that directly answer the query with red circles.
@@
2. We conduct extensive experiments with various popular multi-modal retrievers based on our benchmark, whose results highlight the challenges of LVMR and limitations for existing methods.
@@ 2.2 Multi-modal Embedding Models
Furthermore, with the rise of large language models, some works attempt to transform multi-modal large models (MLLMs) into universal embedders to construct more powerful representations, such as E5-V Jiang et al. (2024) and VLM2VEC Jiang et al. (2025). However, in the video domain, a robust universal embedder has yet to emerge. For this reason, we propose V-Embedder, which outperforms existing video embedders across multiple tasks through synthetic-data-driven training.
@@ 5.1 Annotation Overflow
Table 2: Results on the MomentSeeker benchmark. We report R@1 scores averaged across meta-tasks, with all baselines evaluated in a zero-shot setting. For image-based models (MM-Ret, MagicLens, UniIR, VLM2VEC, E5V), we use the middle frame of videos as input; other models follow their original optimal settings. Additionally, we provide mAP@5 scores and detailed performance on individual sub-tasks in Appendix 8.1.
Method Backbone # Params Frames Meta-Task Average Score Overall
@@
number of tasks - - - 5 5 4 4 18
CLIP-style Visual Encoder
MM-Ret Zhou et al. (2024) CLIP-Base 149M 1 23.2 15.4 10.5 10.5 14.9
MagicLens Zhang et al. (2024) CLIP-Large 428M 1 9.0 2.4 3.2 2.8 4.4
UniIR Wei et al. (2025) CLIP-Large 428M 1 25.0 15.2 6.4 0.0 10.9
LanguageBind Zhu et al. (2023) CLIP-Large 428M 8 39.6 16.4 3.2 0.0 14.8
InternVideo2 Wang et al. (2025) ViT 1B 8 44.6 18.2 4.8 0.0 16.9
@@
CoVR Ventura et al. (2024) BLIP-Large 588M 15 25.8 17.4 12.3 12.3 17.1
@@
VLM2VEC Jiang et al. (2025) LLaVA-1.6 8.4B 8 6.4 6.2 3.0 3.0 4.7
E5V Jiang et al. (2024) LLaVA-1.6 8.4B 8 25.8 16.8 6.2 5.2 13.5
V-Embedder InternVideo2-Chat 8B 8 42.2 20.4 15.0 15.8 23.3
@@ 5.2.1 Baselines
MLLM-based Multi-Modality Encoders. Methods like E5V Jiang et al. (2024) and VLM2VEC Jiang et al. (2025) fall into this category. They utilize large-scale pre-trained multi-modal models to handle flexible modality compositions (e.g., video + text or text alone) and generating multi-modal embeddings.
@@ 5.2.2 Implementation Details
Following previous works Jiang et al. (2025, 2024), we incorporated instructions (e.g., “Represent the given question in one word.” and “Represent the video in one word.”) to guide the model in generating informative embeddings for queries and candidates. All baselines are reproduced using their official implementations with default settings. We evaluate the video models (LanguageBind and InternVideo2) using an input of uniformly sampled 8 frames, while COVR follows its default setting of 15 frames. For image models (E5V, VLM2Vec, UniIR and MagicLens), we use the temporally middle frame as the video input. Instruction usage remains consistent with each baseline’s original configuration. For image-based baselines, we use the middle frame of the video as input. CoVR Ventura et al. (2024) follows its original configuration with 15 frames at a resolution of 384, while LanguageBind, InternVideo2 Wang et al. (2025), and V-Embedder each use 8 frames at a resolution of 224.
@@ 5.3 Main Results
1. All baseline models exhibit suboptimal performance on our benchmark, underscoring the challenge of the tasks. Even state-of-the-art video retrievers like InternVideo2, which achieves 57.9 Recall@1 on Didemo, show drastically reduced effectiveness for long-video temporal grounding (16.9 mean score). This stems from two critical challenges: (1) Extended videos contain numerous hard negative segments requiring fine-grained visual-semantic discrimination, and (2) Instructional queries demand deeper temporal reasoning compared to descriptive queries in standard benchmarks. Current video-text alignment methods prove inadequate for these human-centric instructional scenarios.
2. SOTA Models Exhibit Divergent Strengths and Weaknesses Across Tasks.
  • CLIP-Style Encoders excel in text-based retrieval but falter in multi-modal reasoning. For instance, InternVideo2 achieves strong cross-modal alignment (44.6 on CA, 18.2 on MS), yet performs poorly on tasks requiring explicit multi-modal interaction, scoring only 4.8 on IMS and 0.0 on VMS. This suggests its temporal modeling capabilities are constrained by isolated modality processing.
  • Composed Video Retrievers (e.g., CoVR) invert this trend. Explicit multi-modal fusion enables CoVR to achieve 12.3 VMS (vs. 0.0 for CLIP-style baselines), but sacrifices text retrieval robustness, scoring 25.8 on CA versus InternVideo2’s 44.6. This highlights a trade-off between multi-modal interaction and cross-modal alignment.
  • Composed Image Retrievers (e.g., MM-Ret) demonstrate unexpected generalization to video tasks. Despite relying on single-frame inputs, MM-Ret scores 10.5 on both IMS and VMS, outperforming specialized video encoders like LanguageBind (3.2 IMS). However, its lack of temporal understanding limits overall performance.
  • MLLM-Based Embedding Models (e.g., E5V) dominate instruction-based tasks. E5V surpasses LanguageBind on MS (16.8 vs. 16.4) and VMS (5.2 vs. 0.0), leveraging instruction-tuning for complex queries. Yet, it underperforms in fine-grained temporal alignment (22.3 CA vs. LanguageBind’s 39.6) due to the lack of temporal awareness.
3. Our proposed V-Embedder achieves top performance across all tasks and serves as a strong baseline for future work. V-Embedder unifies temporal perception, multimodal reasoning, and instruction understanding into a unified multimodal video retriever, outperforming the previous SOTA baseline CoVR by 6.2 points. This improvement leverages our backbone model’s robust video understanding capability, inherited from the pretrained video LLM, combined with fine-tuning on our curated multimodal retrieval dataset—together enabling a breakthrough in multimodal comprehension.